OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 4x2 + 9y2 = 36.

OpenStudy (anonymous):

@salehhamadeh

terenzreignz (terenzreignz):

Can you get it to look like this: \[\Large \frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]

OpenStudy (anonymous):

these are the answer choices btw.

OpenStudy (anonymous):

You need to change it to the form that @terenzreignz just mentioned. To do that, you divide both sides by 36 and then complete the squares of both the x and y. The form you get will be: \[(x-x_0)^2/a + (y-y_0)^2/b = 1\]

OpenStudy (anonymous):

ok let me try it!

OpenStudy (anonymous):

The center will be (x0, y0)

OpenStudy (anonymous):

I don't remember how to find the vertices and foci, but you can find the formulas easily in any algebra 2 or precalculus book. Once you get it into the form I mentioned above, all you need to do is plug the numbers into the formulas.

OpenStudy (anonymous):

ok I'm having a bit of trouble getting it into that form

OpenStudy (anonymous):

Where are you getting stuck?

OpenStudy (anonymous):

1x^2/9+1y^2/4=1?

terenzreignz (terenzreignz):

That's good. You can do away with those 1's... \[\Large \frac{x^2}{9}+\frac{y^2}4=1\]

OpenStudy (anonymous):

ok cool! now what do we do?

OpenStudy (anonymous):

@terenzreignz @salehhamadeh

OpenStudy (anonymous):

@terenzreignz 's equation is in the form I gave you. Can't you see the values of x_0, y_0, a, and b?

OpenStudy (anonymous):

x0 and y0 are equal to zero. ie: \[x^2=(x-0)^2\]

OpenStudy (anonymous):

and a and b are 9 and 4, respectively.

OpenStudy (anonymous):

yeah I got that one :)

OpenStudy (anonymous):

@alibea Sorry, the values are +/- 3 and +/- 2, respectively. They should be a^2 and b^2

OpenStudy (anonymous):

Use the formulas on this page to determine what you're looking for: http://en.wikipedia.org/wiki/Ellipse#In_Euclidean_geometry

OpenStudy (anonymous):

ahhhh ok I understand so i think it would be like 0,-3 0,3 right?

OpenStudy (anonymous):

so for focus i got sqrt5 but woul that be sqrt5,0 or 0,sqrt5

OpenStudy (anonymous):

positioning confuses me most of the time

OpenStudy (anonymous):

You need to identify which axis is the major axis and which one is the minor axis. The major axis is the one where the ellipse appears to be more stretched, and vice versa. The foci lie sqrt(5) on the major axis.

OpenStudy (anonymous):

I mean sqrt(5) units from the center on the major axis if this makes more sense.

OpenStudy (anonymous):

if a > b, the x-axis is your major axis and y-axis is your minor axis. If a < b, the y-axis is your major axis. If a=b, then both axes are equal and your equation is an equation for a circle.

OpenStudy (anonymous):

so in this case 9 is greater so the y is greater making the correct choice in this case C right?

OpenStudy (anonymous):

@salehhamadeh @terenzreignz

OpenStudy (anonymous):

a = 9. If a > b, x-axis is the major axis.

OpenStudy (anonymous):

See the graph on this page http://www.mathopenref.com/coordgeneralellipse.html

OpenStudy (anonymous):

yeah but 9 is under y^2 not x^2???

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

9 is b

OpenStudy (anonymous):

OHHHH YOU"RE SO RIGHT!!! nevermind!!!!! :) I'm sorry I see it now. THANK YOU!

OpenStudy (anonymous):

answers actually A lol

OpenStudy (anonymous):

Sorry I had to go and I forgot you were still here. Anyways, I think you get it now. Keep practicing.